Predict the rate of change in the concentration of H2O2 (Δ[H2O2]/Δt).

n the first 11.0 s of the reaction, the concentration of I− drops from 1.000 M to 0.788 M
1
Expert's answer
2018-02-13T11:15:08-0500
A balanced equation of the above mentioned reaction is:
H2O2+3I- +2H+ = (I3)-+2H2O.
The relative reaction rate by iodide-ion is:
Δ[I-]/Δt = (0.788 M -1.000 M)/11 s = -0.01927 M/s.
The relative rate of reaction by hydrogen peroxide is:
-Δ[H2O2]/Δt =-1/3 Δ[I-]/Δt,
Δ[H2O2]/Δt =1/3 Δ[I-]/Δt,
Δ[H2O2]/Δt = 1/3 (-0.01927 M/s) = -6.42·10-3 M/s.
Answer: -6.42·10-3 M/s.